On graph paper, graph the two lines associated with the equations in this system of equations.
3x - 3y = 12
4y - 8x = -16
Determine the solution to the system and check your solution in both equations. Input the solution to the system, putting the solution in ordered pair (x, y) form.

Answer by Alan3354(69443)

On graph paper, graph the two lines associated with the equations in this system of equations.
3x - 3y = 12
4y - 8x = -16
Determine the solution to the system and check your solution in both equations. Input the solution to the system, putting the solution in ordered pair (x, y) form.
----------------
3x - 3y = 12 --> x - y = 4
y = x - 4
---
Pick 2 values for x and find y.
Plot the 2 points and draw a line thru them.
===========================================
4y - 8x = -16 --> y = 2x - 4
----
Do the same for this one.
----
The solution of the system is the point of intersection.